OpenStudy (anonymous):

Write an equation for the parabola whose vertex is at (3,3) and which passes through (5,27).

OpenStudy (anonymous):

General Equation: \[p(x) = a(x-h)^2 + k\] the vertex is the point (h, k). let's plug that in:\[p(x) = a(x-3)^2 + 3\] all we need now is a. you're told that the point (5, 27) exists, let's plug that in: \[27 = a(5-3)^2 + 3\] solve for a. you're parabola general equation would be: \[p(x) = A(x-3)^2 + 3\] where A is whatever "a" is [an actual number]

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (anonymous):

glad i could help :) let me know if you want me to double check your final answer